OpenStudy (anonymous):

Solve for Y Y/3-3=12 simplify your answer as much as possible

OpenStudy (anonymous):

just explain to me how to do it

OpenStudy (saylilbaby):

y=45

OpenStudy (anonymous):

y/3 - 3 = 12 --- add 3 to both sides y/3 = 12 + 3 y/3 = 15 --- now multiply both sides by 3 y = 45 check.. 45/3 - 3 = 12 15 - 3 = 12 12 = 12 (correct) so y = 45

OpenStudy (anonymous):

questions ?

OpenStudy (anonymous):

no thx u